OpenStudy (anonymous):

My worst nihtmare solve 1+tan^2theta= sec^2theta solve for sec^2theta and tan^2theta

zepdrix (zepdrix):

Your equation is already solved for sec^2theta with the way you have it written: \[\large \sec^2\theta=1+\tan^2\theta\] To solve for tan^2theta, subtract 1 from each side,\[\large \tan^2\theta=\sec^2\theta-1\] Is that what you were trying to do? :S

OpenStudy (anonymous):

Well im not too sure but, still thanks zepdrix